# Clustering-based Image-Text Graph Matching for Domain Generalization

### v2

@@ Abstract
Learning domain-invariant visual representations is important to train a model that can generalize well to unseen target task domains. Recent works demonstrate that text descriptions contain high-level class-discriminative information and such auxiliary semantic cues can be used as effective pivot embedding for domain generalization problem. However, they use pivot embedding in global manner (i.e., aligning an image embedding with sentence-level text embedding), not fully utilizing the semantic cues of given text description. In this work, we advocate for the use of local alignment between image regions and corresponding textual descriptions. To this end, we first represent image and text inputs with graphs. We subsequently cluster nodes in those graphs and match the graph-based image node features into textual graphs. This matching process is conducted globally and locally, tightly aligning visual and textual semantic sub-structures. We experiment with large-scale public datasets, such as CUB-DG and DomainBed, and our model achieves matched or better state-of-the-art performance on these datasets. Our code will be publicly available upon publication.
@@ 1 Introduction
Domain generalization aims to improve a model’s generalization ability for unseen task domains. Previous research has explored various approaches to address this challenge, including minimizing domain discrepancies in the visual feature space [32, 45, 15, 24], augmenting data to cover various domains [52, 51, 50, 23], and utilizing ensemble learning [10, 28]. Notably, recent work such as GVRT [36] suggests leveraging natural language descriptions (e.g., “this bird has a blue crown, green primaries, and a yellow belly”) to infuse visual encoders with domain-invariant semantic cues, i.e., a visual encoder is optimized to produce an embedding that aligns well with the corresponding text embedding. While promising, optimizing a model with such global alignment often leads to suboptimal results, as these models may lack diverse attribute focus and occasionally attend to irrelevant regions for the class (e.g., see Figure 5(a)).
@@ 2 Related Work
Domain generalization aims to enhance a model’s ability to generalize to unseen target domains with different data distributions compared to the source domains. The main idea of domain generalization is to learn domain-invariant features from multiple source domains. Various methods have been proposed to resolve this problem (i) by reducing domain discrepancies in the feature space [32, 45, 15, 24], (ii) by implementing data augmentation [52, 51, 50, 23], and (iii) by utilizing ensemble learning [10]. (iv) Other studies have proposed using auxiliary semantic cues to facilitate learning domain-invariant features [3, 42, 11].
@@ 4.1 Setup
Same as previous domain generalization approaches, we also use ResNet-50 [20], pre-trained on ImageNet [13], as our backbone, yielding a 2,048-dimensional visual representation from the last layer. Our model is trained end-to-end for 5,000 training steps using Adam optimizer with a learning rate of 5e-5. For training, we use standard image augmentation techniques such as random cropping, horizontal flipping, color jittering, grayscale conversion, and normalization. Our implementation is based on DomainBed [17], which is a unified domain generalization testbed, and our code will be publicly available upon publication. More details, including information that varies depending on the dataset, are available in supplementary material.
@@ 4.2 Performance Comparison
(0.5$\uparrow$) (0.7$\uparrow$) (0.9$\uparrow$) (0.5$\uparrow$) (0.7$\uparrow$)
As shown in Table 1, we compared the out-of-distribution classification accuracies on the following two datasets: CUB-DG (top) and PACS (bottom) datasets. We compare ours with other existing state-of-the-art domain generalization approaches, SagNet [37], MIRO [11], SD [39], CORAL [45], GVRT [36], SelfReg [24], and mDSDI [7]. Due to space constraints, we only report top-5 results (we provide full tables in supplementary material).
As shown in Table 1 (top), our proposed method clearly outperforms the other domain generalization techniques on the CUB-DG dataset in all target domains with a significant gain. In terms of the average image classification accuracy, ours shows 59.1%, which is 2.1% higher than GVRT [36] (which uses the same image and text inputs) and 3.7%-14.3% higher than other image-only approaches. Similar trends are also observed in our experiment on the large-scale PACS [30] dataset. As shown in Table 1 (bottom), our model also outperforms the other approaches, i.e., ours shows 87.0% that is 0.7% higher than SagNet [37]-based SOTA approach and 1.9% higher than GVRT [36]. These confirm that our graph-based approach is effective in aligning visual and textual encoders for fine-grained image classification tasks, improving the visual encoder’s generalization power to unseen target domains.
@@ 4.4 Analysis on CUB-DG Dataset
As shown in Figure 4, we provide t-SNE [35] visualization of (a) ERM [47], (b) GVRT [36], and (c) Ours to visualize their embedding space on CUB-DG dataset. We use different marker styles (for target domains) and different colors (for classes). An ideal model would show that visual features of the same class but different domains are gathered together. Ours clearly outperform ERM, which has scattered points per domain and better than GVRT in that features of the same class but different domains are more clustered (see red boxes). In Figure 4 (d), we provide box plots for GVRT and ours, showing that our model produces lower same-class inter-domain distances than GVRT. Note that we provide detailed t-SNE visualizations in supplementary material.
@@ 4.5 Performance on DomainBed Benchmark
We evaluate our model with a large-scale DomainBed [17] datasets. We use the following five multi-domain datasets, including VLCS [14], PACS [30], OfficeHome [48], TerraIncognita [4], and DomainNet [38], comparing ours with 19 domain generalization algorithms. Due to space constraints, we only report top-5 results (see supplementary material for full table). The reported score represents averaged results obtained from three independent runs using randomly chosen hyperparameters. We observe in Table 4 that our proposed method shows matched or better state-of-the-art performance, where it ranks 1st (tied) in average performance.

### v3

@@ Abstract
Learning domain-invariant visual representations is important to train a model that can generalize well to unseen target task domains. Recent works demonstrate that text descriptions contain high-level class-discriminative information and such auxiliary semantic cues can be used as effective pivot embedding for domain generalization problems. However, they use pivot embedding in a global manner (i.e., aligning an image embedding with sentence-level text embedding), which does not fully utilize the semantic cues of given text description. In this work, we advocate for the use of local alignment between image regions and corresponding textual descriptions to get domain-invariant features. To this end, we first represent image and text inputs as graphs. We then cluster nodes within these graphs and match the graph-based image node features to the nodes of textual graphs. This matching process is conducted both globally and locally, tightly aligning visual and textual semantic sub-structures. We experiment with large-scale public datasets, such as CUB-DG and DomainBed, and our model achieves matched or better state-of-the-art performance on these datasets. The code is available at: https://github.com/noparkee/Graph-Clustering-based-DG
@@ 1 Introduction
Domain generalization aims to improve a model’s generalization ability for unseen task domains. Previous research has explored various approaches to address this challenge, including minimizing domain discrepancies in the visual feature space [45, 23], augmenting data to cover various domains [50, 22], and utilizing ensemble learning [10, 27]. Notably, recent work such as GVRT [36] suggests leveraging natural language descriptions (e.g., “this bird has a blue crown, green primaries, and a yellow belly”) to infuse visual encoders with domain-invariant semantic cues, i.e., a visual encoder is optimized to produce an embedding that aligns well with the corresponding text embedding. While promising, optimizing a model with such global alignment often leads to suboptimal results, as these models may lack diverse attribute focus and occasionally attend to irrelevant regions for the class (e.g., see Figure 6(a)).
@@ 2 Related Work
Domain generalization aims to enhance a model’s ability to generalize to unseen target domains with different data distributions compared to the source domains. The main idea of domain generalization is to learn domain-invariant features from multiple source domains. Various methods have been proposed to resolve this problem (i) by reducing domain discrepancies in the feature space [45, 23], (ii) by implementing data augmentation [50, 22], and (iii) by utilizing ensemble learning [10]. (iv) Other studies have proposed using auxiliary semantic cues to facilitate learning domain-invariant features [3, 42, 11].
@@ 4.1 Setup
Same as previous domain generalization approaches, we also use ResNet50 [19], pre-trained on ImageNet [13], as our backbone, yielding a 2,048-dimensional visual representation from the last layer. Our model is trained end-to-end for 5,000 training steps using Adam optimizer with a learning rate of 5e-5. For training, we use standard image augmentation techniques such as random cropping, horizontal flipping, color jittering, grayscale conversion, and normalization. Our implementation is based on DomainBed [17], which is a unified domain generalization testbed, and our code will be publicly available upon publication. More details, including information that varies depending on the dataset, are available in supplementary material.
@@ 4.2 Performance Comparison
As shown in Table 1, we compared the out-of-distribution classification accuracies on the following two datasets: CUB-DG (top) and PACS (bottom) datasets. We compare ours with other existing state-of-the-art domain generalization approaches, SagNet [37], MIRO [11], SD [39], CORAL [45], GVRT [36], SelfReg [23], mDSDI [7], and CCFP [29]. Due to space constraints, we only report top-6 results (we provide full tables in supplementary material).
As shown in Table 1 (top), our proposed method clearly outperforms the other domain generalization techniques on the CUB-DG dataset in all target domains with a significant gain. In terms of the average image classification accuracy, ours shows 59.1%, which is 2.1% higher than GVRT [36] (which uses the same image and text inputs) and 3.1% higher than image-only approach, CCFP [29]. Similar trends are also observed in our experiment on the large-scale PACS [30] dataset. As shown in Table 1 (bottom), our model also outperforms the other approaches, i.e., ours shows 87.0% that is 0.4% higher than CCFP [29]-based SOTA approach and 1.9% higher than GVRT [36]. These confirm that our graph-based approach is effective in aligning visual and textual encoders for fine-grained image classification tasks, improving the visual encoder’s generalization power to unseen target domains.
@@ 4.4 Analysis on CUB-DG Dataset
As shown in Figure 4, we provide t-SNE [35] visualization of
Figure 5: Exemplars of the nearest examples from PACS dataset (in the unseen target domain) to the given image (e.g., “dog” and “horse”).
(a) ERM [47], (b) GVRT [36], and (c) Ours to visualize their embedding space on CUB-DG dataset. We use different marker styles (for target domains) and different colors (for classes). An ideal model would show that visual features of the same class but different domains are gathered together. Ours clearly outperform ERM, which has scattered points per domain and better than GVRT in that features of the same class but different domains are more clustered (see red boxes). In Figure 4 (d), we provide box plots for GVRT and ours, showing that our model produces lower same-class inter-domain distances than GVRT. Note that we provide detailed t-SNE visualizations in supplementary material. Figure 5 further shows the nearest examples in the unseen target domain to the given image (e.g., “dog”). In contrast to MIRO and GVRT, which often provide examples of different classes (e.g., “person”), ours consistently provide examples of the same class. This is consistent with our t-SNE analysis. We provide more examples in in supplementary material.
@@
Table 5: Performance comparison between variants of our model with different matching techniques: bipartite matching and greedy matching. Data: CUB-DG.
Bipartite Matching Greedy Matching Time Complexity Target Domain Avg.
Photo Cartoon Art Paint
- $O(V^{2})$ 75.3 64.7 54.8 36.6 57.8
- $O(V^{3})$ 75.4 65.5 54.0 41.4 59.1
4.4.5 Complexity of Graph Matching.
In our matching algorithm design, we explored two approaches: bipartite matching and greedy matching. Bipartite matching establishes one-to-one correspondences between clusters, minimizing pairwise distances, while greedy matching allows many-to-one associations based on spatial proximity. Bipartite matching operates with a time complexity of $O(V^{3})$, where $V$ denotes the number of vertices, while greedy matching operates in $O(V^{2})$. Despite the higher time complexity of bipartite matching, our experimental results (refer to Table 5) demonstrate its superior performance over greedy matching. This may be attributed to the constraints imposed on one-to-one matching, which result in a dispersed effect on models attempting to optimize cluster pairs on a global scale. Moreover, given that $V$ does not exceed 5 in our method, we opted for bipartite matching due to its enhanced performance in our specific context.
@@ 4.5 Performance on DomainBed Benchmark
We evaluate our model with a large-scale DomainBed [17] datasets. We use the following five multi-domain datasets, including VLCS [14], PACS [30], OfficeHome [48], TerraIncognita [4], and DomainNet [38], comparing ours with 19 domain generalization algorithms. Due to space constraints, we only report top-6 results (see supplementary material for full table). The reported score represents averaged results obtained from three independent runs using randomly chosen hyperparameters. We observe in Table 4 that our proposed method shows matched or better state-of-the-art performance, where it ranks 1st (tied) in average performance.